OpenStudy (anonymous):

Algebra Two Help: Given the function f(x) = log5(x + 1), find the value of f^−1(2).

OpenStudy (lucaz):

find the value of what?

OpenStudy (lucaz):

inverse ?

OpenStudy (anonymous):

inverse of the function f(x)=log5(x+1)

OpenStudy (lucaz):

y = log5(x+1) you hav to write this in the exponential form from the function you see that the base 10 raised to the power of y is equal to 5(x+1) right?

OpenStudy (anonymous):

The base is five. I'm not sure how to properly type out logs. The exponential form would be x=5^(x+1), right?

OpenStudy (lucaz):

oh, so it's log base 5 of (x+1)?

OpenStudy (anonymous):

yes

OpenStudy (lucaz):

right, then the base 5^y = x+1, in exponential form

OpenStudy (lucaz):

subtracting 1 from both sides 5^y -1 = x we're looking for the inverse, then we swap x and y 5^x - 1 = y f^-1(2) = 5^(2) - 1

OpenStudy (anonymous):

so 24

OpenStudy (lucaz):

yeah

OpenStudy (anonymous):

Thank you so much1

OpenStudy (lucaz):

no problem